OpenStudy (kamille):

Hello, I need to solve for x. \[2\sin ^{2}x=3\sin(\frac{ 3\pi }{ 2 }+x)\] Firstly, do I need to reduction 3\sin(\frac{ 3\pi }{ 2 }+x)\]?

OpenStudy (kamille):

\[2\sin ^{2}x=3\sin(\frac{ 3\pi }{ 2 }+x)\]

OpenStudy (kamille):

@InYourHead

OpenStudy (anonymous):

I think you can reduce this first : \[\sin(\frac{3 \pi}{2} + x) = ??\]

OpenStudy (kamille):

cosx?

OpenStudy (anonymous):

How?? Any logic??

terenzreignz (terenzreignz):

I'd love for that to be true, but unfortunately, the truth hurts XD \[\huge \sin(\alpha+\beta)=\sin\alpha\cos\beta+\cos\alpha\sin\beta\]

OpenStudy (anonymous):

cos(x)?? Something left I think, check it further..

OpenStudy (kamille):

Because when there is 3pi/3 fuction chages a "name". So it becomes from sin to cos, right?

OpenStudy (anonymous):

@terenzreignz I think there is no need to go by that formula..

OpenStudy (kamille):



OpenStudy (kamille):

and i have forgotten -, right?

OpenStudy (kamille):

correction 3pi/2!

terenzreignz (terenzreignz):

Don't confuse it with \[\huge \sin\left(\frac{\pi}{2}-\theta\right)=\cos \ \theta\]

OpenStudy (anonymous):

is x acute?

OpenStudy (anonymous):

Yep,, So it will be simply -cos(x) Getting??